Solve, and express your answer in exact form 3^x = 4^(l-x)
Answer by ankor@dixie-net.com(12692)

Solve, and express your answer in exact form 3^x = 4^(l-x)
:
Not sure what exact form is, but here is a method using natural logs.
:
ln(3^x) = ln[4^(1-x)]
:
Use the log equiv of exponents:
x*ln(3) = (1-x)*ln(4)
:
Find the logs of 3 & 4:
1.0986x = 1.38629(1-x)
1.0986x = 1.38629 - 1.38629x
1.0986x + 1.38629x = 1.38629
2.48489x = 1.38629
x = 1.38629/2.48489
x = .55789 (rounded)
:
:
Check on a good calc:
enter ln(3^.55789) = .61289
enter ln(4^.44211) = .61290; our solution should be exact enough